OpenStudy (anonymous):

Show the expanded form and simplification of x4/x2. Explain in your own words how you can simplify x4/x2 without having to write the expanded form.

OpenStudy (anonymous):

all the numbers are exponets

OpenStudy (anonymous):

The expanded form should be: \[\frac{ x*x*x*x }{ x*x }\] The simplified form should be (if you cross two from the top and two from the bottom): \[x*x=x^2\] You can simplify x4/x2 with the index law: \[\frac{ a^m }{ a^n }=a^{m-n}\] which can be used for: \[\frac{ x^4 }{ x^2 }=x^{4-2}=x^2\]

OpenStudy (anonymous):

Create your own fraction with like bases, coefficients, and show its simplification